What is the energy of photon of wavelength 1 mm? Compare this with the characteristic thermal energy at 3000K.​


1
Expert's answer
2021-01-22T15:04:58-0500

"E=h\nu=\frac{hc}{\lambda}=\frac{6.626\cdot10^{-34}\cdot3\cdot10^8}{1\cdot10^{-3}}=19.878\cdot10^{-23}(J)"


The characteristic thermal energy


"E_t\approx kT=1.38\cdot10^{-23}\cdot3000=4.14\cdot10^{-20}(J)"


"E_t/E\approx208"


So, the characteristic thermal energy is 208 times greater than the energy of photon. Answer